A particle is moving with 3 times faster than speed of electron. Ratio of wavelength of particle and electron is 1.8*10^-4, then the particle is:-

 

a) Neutron

b) Alpha particle

c) Deutron

d) Tritium

 

Grade:12th pass

1 Answers

Rohan Mathur
5 years ago

A neutron.

Explanation:

We use the de Broglie expression:

λ=hmv

So we can write:

λe=hmeve (1)

λp=hmpvp (2)

vp=3ve (3)

Substituting (3) into (2)

λp=hmp3ve (4)

Dividing (4) by (3)

λpλe=hmp3ve.meveh=1.8×104

me3mp=1.8×104

memp=5.4×104

This ratio is 5.39522×104 for a neutron and 5.36099×104 for a proton, so a neutron seems to be the most likely candidate for the particle in question, though I guess there could be others.